# B-cos LM: Efficiently Transforming Pre-trained Language Models for Improved Explainability

### v2

@@ 7 B-cosifying Decoder-Only Models for Generation Tasks
Table 7: Alignment results (MRR) on BLiMP and IOI. BLiMP MRR scores are averaged over nine subsets, and Logit and ALTI Logit results are replicated from the original paper (Ferrando et al., 2023). Best scores are marked in bold. B-cos explanations achieve the best alignment with ground truth evidence. Results for each subset can be found in Table 14 and Table 15 in Appendix N.
Results
@@ 8 Conclusion
In this work, we introduce B-cos LM, a bias-free dynamic linear model that learns task-relevant patterns through increased input-weight alignment pressure. B-cos LMs generate more faithful and human interpretable explanations while maintaining strong task performance and fast convergence. Based on our in-depth analysis of B-cosification, we provide three recommendations for effectively transforming PLMs into B-cos LMs: (1) combine B-cos conversion and task fine-tuning for efficient B-cosification. If resources allow, additional B-cos pre-training can further improve task performance and explanation faithfulness; (2) carefully select the parameter B, as excessively large values can reduce model capacity and lead to overly sparse explanations; and (3) be mindful of biases in training data, especially at high B values, as B-cosification may amplify existing biases.
We also explore adapting decoder-only models into B-cos LMs for generation tasks and show that, with additional training, they match the language modeling performance of conventional models while providing better explanations. We hope these findings support future efforts in building explainable LLMs.
@@ 9 Limitations
This study has certain limitations that should be acknowledged.
Firstly, the automatic evaluation metrics we use may not fully capture the faithfulness of different explanation methods (Feng et al., 2018; Lapuschkin et al., 2019). However, since there is no universal consensus on the most reliable evaluation metrics, this remains an open challenge in explainability research.
@@
10 Broader Impact Statement
As discussed in § 6, B-cos LMs can overfit to biases present in the training data. Although their more faithful and human interpretable explanations make biased predictions easier to detect, this does not eliminate the risk of unintended bias amplification. We encourage users to carefully assess potential biases in their specific use cases before deploying B-cos LMs and to incorporate bias mitigation strategies where necessary.
@@ Appendix B Notation
Appendix C Implementation Details
@@
Appendix D SeqPG Example
@@
Appendix E Task Performance of Other B-cos LMs
@@
Appendix F Faithfulness Evaluation of Other B-cos LMs
@@
Appendix G Comparison to Rationale-Based Models
@@
Appendix H Human Evaluation Details
@@
Appendix I More Examples of B-cos Explanations
@@
Appendix J Impact of B on Input-weight Alignment
@@
Appendix K Effects of B on Other Metrics
@@
Appendix L B-cos Explanations with Different B Values
@@
Appendix M Example of Model Bias
@@
Appendix N Decoder-Only Model Results

### v3

@@ 8 Practical Guidance for Training B-cos LMs
Table 6: Alignment results (MRR) on BLiMP and IOI. Logit and ALTI Logit results are replicated from the original paper (Ferrando et al., 2023). Best scores are marked in bold. B-cos explanations achieve the best alignment with ground truth evidence. Results for each subset can be found in Table 14 and Table 15 in Appendix P.
@@ 9 Conclusion
In this work, we introduce B-cos LM, a bias-free dynamic linear model that learns task-relevant patterns through increased input-weight alignment pressure. B-cos LMs generate more faithful and human interpretable explanations while maintaining strong task performance and fast convergence. We also explore adapting decoder-only models into B-cos LMs for generation tasks and show that, with additional training, they match the language modeling performance of conventional models while providing better explanations. Based on our in-depth analysis of B-cosification, we also provide a practical guideline for effectively transforming PLMs into B-cos LMs.
@@ 10 Limitations
This study has certain limitations that should be acknowledged. First, the automatic evaluation metrics we use may not fully capture the faithfulness of different explanation methods (Feng et al., 2018; Lapuschkin et al., 2019). However, since there is no universal consensus on the most reliable evaluation metrics, this remains an open challenge in explainability research.
@@
Finally, B-cos explanations, like all input-based approaches, are not suitable for all tasks. They are particularly suitable for NLP tasks where predictions can be directly attributed to specific words or phrases in the input. Examples include various text classification tasks, such as sentiment analysis and toxicity detection, where the presence or absence of certain tokens often provides clear evidence for the model’s decision. In such settings, token-level explanations are both intuitive and precise, enabling users to easily verify whether the model attends to the expected parts of the input. On the contrary, in tasks that involve multi-step reasoning or external knowledge retrieval such as commonsense reasoning or fact verification, token-level explanations are often less informative. In these cases, higher-level explanations, such as concept-based or natural language explanations, may better capture the underlying reasoning process and provide users with more meaningful insights.
Acknowledgements
This work was funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – GRK 2853/1 “Neuroexplicit Models of Language, Vision, and Action” - project number 471607914. We also thank all those who provided valuable feedback and suggestions during the development of this work. Their input has been instrumental in improving its quality.
@@ Appendix B Notation
Appendix C Dynamic Linear Representation of Model Components
B-cos Layers
B-cos layers are designed as bias-free, dynamic linear modules with a dynamic linear weight matrix $\mathbf{W(x)=|\text{cos}(x,\hat{W})|^{\text{B-1}}\otimes\hat{W}}$. Here, $\mathbf{\otimes}$ scales the rows of the matrix $\mathbf{\hat{W}}$ to its right by the scalar entries of the vector to its left.
Non-linear Activation Functions
Attention Blocks
$\text{MSA}\mathbf{(X)=U[W}_{1}\mathbf{(X)X,...,W}_{H}\mathbf{(X)X)]}$ (5)
Normalization Layers
where $\mathbf{\mathcal{X}}$ represents a batch or sequence of inputs and $\star$ is the dimension along which the mean $\langle\cdot\rangle$ and variance $\text{var}(\cdot)$ are computed (e.g., across the batch or layer). Unlike standard normalization, this variant omits the bias term in the affine transformation to preserve explanation completeness. If a running mean estimate is used during inference, the centering term $\mathbf{\langle\mathcal{X}\rangle_{\star}}$ is also removed. This yields a bias-free, dynamic linear transformation with weight $\sqrt{\text{var}\mathbf{{}^{-1}_{\star}(\mathcal{X})}}\mathbf{\times\gamma}$.
Appendix D Implementation Details
@@
Appendix E Explanation Efficiency
Method Time (s) Memory (GB)
ShapSampl 37.22 21.95
LIME 6.82 21.96
SIG 67.46 29.09
DecompX 0.76 48.38
B-cos 0.08 2.82
Appendix F SeqPG Example
@@
Appendix G Task Performance of Other B-cos LMs
@@
Appendix H Faithfulness Evaluation of Other B-cos LMs
@@
Appendix I Comparison to Rationale-Based Models
@@
Appendix J Human Evaluation Details
@@
Appendix K More Examples of B-cos Explanations
@@
Figure 13 presents two examples where B-cos explanations do not align with human expectations. As B-cos explanations faithfully summarize model computations, they may also reveal spurious correlations from the training data and occasionally produce unreasonable rationales. For example, in the first case, the token “[user]” receives high attribution, likely due to recurring patterns in the data. Another observed error is polarity misinterpretation: in the second case, the token “niggas” is incorrectly highlighted as positive evidence for a non-toxic prediction. Similar types of undesired behaviors have also been reported for other explanation methods.
@@
Figure 13: Two undesired B-cos explanations from B-cos BERT on HateXplain. Green (red) indicates the positive (negative) impact of tokens on the prediction. B-cos explanations can be influenced by spurious correlations learned by the model and may occasionally misattribute the contribution of certain tokens.
Appendix L Impact of B on Input-weight Alignment
@@
Appendix M Effects of B on Other Metrics
@@
Appendix N B-cos Explanations with Different B Values
@@
Appendix O Example of Model Bias
@@
Appendix P Decoder-Only Model Results